An ion with m/z of 100 is born at a distance 5mm in front of a repeller plate with an applied voltage of -1500V. An extraction grid is 2cm away from the repeller with an applied potential of +500V. What velocity does the ion experience exiting the grid? What would happen if the ion had an initial velocity but in the backward direction (toward the repeller plate)?
Answer #1

Mass to charge ratio m / z is 100

Voltage of repeller = - 1500 v

Voltage of grid = 500 v

Voltage difference = 2000 volt

Distance between repeller and grid = 2 cm = 0.02 m

Distance of ion from repeller = 5 mm = 0.005 m

Assuming uniform electric field between both the plates

Energy gained by ion = 2000×z× (.195/0.2) = 1950z. Joule

Energy will be get converted to kinetic energy

Hence. 0.5×m×V2 = 1950z

Or. V = 6.24 m/s

Hence speed of ion will be 6.24 m/s

If ion had an initial velocity in the direction of repeller , it will experience resistivive force and finally it will turn and come out througj grid.